OpenStudy (anonymous):

Find the roots for the following quadratic equation using the quadratic formula: 3x^2 - 5x + 2 = 0 Leave answers as fractions or write as an integer separated by a comma.

OpenStudy (anonymous):

if you only need a solution try this http://www.wolframalpha.com/input/?i=3x^2+-+5x+%2B+2+%3D+0+

OpenStudy (anonymous):

it will give it to you in a nano second otherwise you have to compute \[\frac{-(-5)\pm\sqrt{(-5)^2-4\times 3\times 2}}{2\times 3}\]

OpenStudy (anonymous):

oh. ok

OpenStudy (anonymous):

wait.. whats the answer

OpenStudy (anonymous):

did you check the link i sent? it is in it